OpenStudy (anonymous):

Solve the initial-value problem: x*(dy/dx)-y/(x+1)=x, y(1)=0, x>0 Please show your work! I'm confused.

OpenStudy (anonymous):

\[xy'-\frac{1}{x+1}y=x\\ y'-\frac{1}{x(x+1)}y=1\] Linear ODE, so find the integrating factor: \[\mu(x)={\large e}^{\displaystyle\int\frac{dx}{x(x+1)}}\]

OpenStudy (anonymous):

Thanks~!

OpenStudy (anonymous):

yw